### v1

# Efficient Artifacts Removal for Adaptive Deep Brain Stimulation and a Temporal Event Localization Analysis

Abstract.
Objective. Adaptive deep brain stimulation (aDBS) leverages symptom-related biomarkers to deliver personalized neuromodulation therapy, with the potential to improve treatment efficacy and reduce power consumption compared to conventional DBS. However, stimulation-induced signal contamination remains a major technical barrier to advance its clinical application. Approach. Existing artifact removal strategies, both front-end and back-end, face trade-offs between artifact suppression and algorithmic flexibility. Among back-end algorithms, Shrinkage and Manifold-based Artifact Removal using Template Adaptation (SMARTA) has shown promising performance in mitigating stimulus artifacts with minimal distortion to local field potentials (LFPs), but its high computational demand and incapability of handling transient direct current (DC) artifacts limits its use in real-time applications. To address this, we developed SMARTA+, a computationally efficient extension of SMARTA capable of suppressing both stimulus artifacts and DC transient artifacts while supporting flexible algorithmic design. Main results. We evaluated SMARTA+ using semi-real aDBS data and real data from Parkinson’s disease patients. Compared to SMARTA and other established artifact suppression methods, SMARTA+ achieved comparable or superior artifact removal while significantly reducing computation time. It preserved the spectral and temporal structure, ranging from beta band to high frequency oscillation, of the underlying LFPs and demonstrated robustness across a variety of simulated stimulation protocols. Furthermore, the temporal event localization analysis show that SMARTA+ can more accurately determine the beta burst events compared with existing tools. Significance. By enhancing artifact suppression and improving computational efficiency, we show that SMARTA+ is potential for advancing real-time, closed-loop aDBS systems for neuromodulation therapies across diverse neurological disorders.
Keywords: Deep brain stimulation; transient DC artifact; stimulus artifact; artifact removal; optimal shrinkage; temporal event localization analysis; SMARTA+
T.-C. Liu: Neuroscience Research Center, Chang Gung Memorial Hospital, Taoyuan, Taiwan; Department of Mathematics, National Taiwan University, Taipei, Taiwan.
P.-L. Chen: Neuroscience Research Center and Division of Movement Disorders, Department of Neurology, Chang Gung Memorial Hospital, Taoyuan, Taiwan.
Y.-C. Chen: Division of Movement Disorders, Department of Neurology, and College of Medicine, Chang Gung University, Taoyuan, Taiwan.
P.-H. Tu: College of Medicine, Chang Gung University, Taoyuan, Taiwan; Department of Neurosurgery, Chang Gung Memorial Hospital at Linkou; School of Medicine, National Tsing Hua University, Hsinchu, Taiwan.
C.-H. Yeh: Department of Neuroradiology, Chang Gung Memorial Hospital, Linkou, Taiwan.
M.-C. Yeap: Department of Neurosurgery, Chang Gung Memorial Hospital at Linkou, Taoyuan, Taiwan.
$C.$ Chen, co-corresponding author: neurozoe@gmail.com; affiliated with Neuroscience Research Center, Division of Movement Disorders, and College of Medicine, Chang Gung University, Taoyuan, Taiwan.
H.-T. Wu, co-corresponding author: hauwu@cims.nyu.edu; Courant Institute of Mathematical Sciences, New York University, New York, USA.
1. Introduction
Deep brain stimulation (DBS) is a widely used and effective therapy for advanced Parkinson’s disease (PD) [21, 3, 20, 8]. By delivering electrical stimulation through implanted electrodes in the subthalamic nucleus (STN), DBS can alleviate motor impairments and improve patients’ quality of life. Additionally, local field potentials (LFPs) can be recorded from the same electrodes during stimulation, providing real-time insights into PD-related neural activity. Biomarkers derived from LFP enable adaptive deep brain stimulation (aDBS), where stimulation parameters are dynamically adjusted to respond to fluctuations in disease state [23, 2]. The clinical benefits of aDBS have been demonstrated in PD, with improvements observed in motor symptoms such as tremor [40] and bradykinesia [22]. Beyond its application in PD, aDBS has also shown promise in treating essential tremor [41, 29] and epilepsy [19, 33]. These studies suggest that aDBS may be broadly applicable to various movement disorders, offering a more efficient and patient-specific therapeutic strategy.
[…]
Various techniques, spanning both front-end and back-end processing, have been developed to mitigate stimulation artifacts in aDBS recordings. A basic front-end approach is common-mode rejection, where two sensing contacts adjacent to the stimulation electrode are used to obtain bipolar LFPs by signal subtraction [23, 2]. Ideally, when contacts are symmetrically positioned around the stimulation site, artifacts cancel out. However, impedance mismatches and tissue asymmetries often leave residual artifacts and DC transient distortions [1]. To further reduce artifacts, blanking techniques can be applied at the front-end [32] or back-end [14], temporarily disabling signal acquisition during stimulation pulses. While effective at suppressing high-amplitude artifacts, blanking also removes underlying neural signals and fails to address transient DC artifacts that persist beyond the stimulation period. When the frequency band of interest is known in advance, such as the beta band in PD therapy [23, 2], careful frequency planning is a commonly used strategy, as previously discussed. However, this approach becomes less effective in the presence of large artifacts and constrains the flexibility of stimulation parameters, thereby limiting the potential for optimizing therapeutic efficacy. Digital filtering and template subtraction (TS) are widely used back-end methods. TS assumes artifact stability over time, averaging repeated instances to subtract from the signal [13, 35]. However, time-varying brain responses may alter artifact characteristics [24], limiting TS performance. To address this, Shrinkage and Manifold-based Artifact Removal using Template Adaptation (SMARTA) was introduced [24]. By leveraging manifold denoising and random matrix theory, SMARTA adapts templates for individual artifacts using the underlying geometry of stimulation-induced distortions and the high-dimensional nature of LFP noise. This enables precise artifact removal while preserving neural signals. Despite its strengths, SMARTA has two key limitations. First, it is computationally intensive: processing a 7-ms segment takes over 200 ms, making it unsuitable for real-time aDBS. Second, like TS, it does not resolve DC transient artifacts, which limits its use for short or repeated stimulation periods and impairs beta burst detection. To support real-time aDBS with an accurate beta burst detection, a more efficient and DC-transient-aware variant of SMARTA is needed for accurate and timely biomarker extraction.
In this study, we refined SMARTA to address its limitations, resulting in SMARTA+, which preserves SMARTA’s core strengths while greatly improving computational efficiency and mitigating DC transient artifacts. A key enhancement involved replacing SMARTA’s most computationally intensive step, the k-nearest neighbors (KNN) search, with an approximate nearest neighbors (ANN) algorithm [4, 17, 18]. Unlike KNN, which performs exhaustive searches, ANN uses decision trees to identify similar artifacts with significantly reduced computational burden [4]. Additionally, we introduced a dimensionality reduction step based on wavelet transforms, which exploits the sparsity structure of stimulus artifacts [9]. These modifications enable the construction of a large stimulus artifact library, potentially from other subjects, allowing SMARTA+ to match each artifact with the most similar examples across datasets. This cross-subject learning improves artifact estimation and suppression, especially during stimulation onset, where DC transients alter artifact morphology (see Figure 1). Without a sufficiently diverse library, onset artifacts are particularly difficult to model. To further suppress DC transients, we incorporated a line-fitting algorithm that estimates and removes the slow-varying baseline shift at the start of each stimulation period. Together, these advances enable SMARTA+ to achieve effective artifact removal with a fraction of the original computational cost, advancing it toward real-time aDBS deployment.
Figure 1. Visualization of stimulus artifacts recorded during aDBS using principal component analysis (PCA), with the top three principal components shown. The color bar indicates the timing of each artifact relative to stimulation onset. Notably, artifact patterns near the onset differ markedly from those at later times, revealing changes in artifact dynamics over time. Since PCA is a linear dimensionality reduction method, the observed structure suggests underlying nonlinear structure and nontrivial dynamics among the artifacts.
We evaluated the performance of SMARTA+ using aDBS signals recorded from PD patients and semi-real aDBS signals synthesized from these recordings. In addition to standard metrics, including normalized mean square error (NMSE) for quantifying LFP recovery across spectral bands, artifact residual (AR) for time-domain suppression, and spectral concentration (SC) for spectral contamination [24], we introduced a temporal event localization analysis to assess how accurately SMARTA+ recovers beta burst events. Specifically, we evaluated recall, precision, and F1-score to quantify the detection of beta burst onset and offset. This temporal event localization analysis, while previously applied in sEMG-based closed-loop DBS systems [30] for evaluating closed-loop DBS with post-stimulation control signals and sleep apnea detection from the photoplethysmogram signal [5], have not yet been used to evaluate aDBS algorithms that rely on real-time beta activity from LFPs. To our knowledge, SMARTA+ is the first algorithm explicitly designed to preserve beta burst timing. For benchmarking, we compared its performance against template subtraction [7], SMARTA [24] and blanking methods [1].
2. Materials
Patients undergoing bilateral implantation of STN-DBS for the treatment of advanced PD were recruited. The surgical procedures were described in [6]. LFPs were recorded one day after electrode implantation through the externalized electrode using Neuro Omega (Alpha Omega Engineering, Israel). All surgeries and recordings were conducted at Chang Gung Memorial Hospital, Linkou, Taoyuan, Taiwan, between 2023 and 2024 by the coauthors Dr. P. H. Tu and Dr. M. C. Yeap.
[…]
For cDBS recordings, stimulation started 10 seconds after session onset and ended 10 seconds before completion. The aDBS protocol, particularly the threshold, is summarized in the next subsection. LFPs were recorded using the Neuro Omega system and processed in real time in MATLAB2019a. Stimulation was triggered if beta amplitude exceeded the threshold and remained on for at least 400 ms; it was deactivated only when the amplitude fell below the threshold. For each patient, aDBS was recorded for 120 s using the same current amplitude as cDBS.
2.1. aDBS protocol
The aDBS protocol followed the method in [23]. A second-order Butterworth band-pass filter (3-37 Hz) was first applied to the baseline LFP signal, the peak beta frequency (13-35 Hz) was identified, and a peak IIR filter with a Q factor of 3 centered at this frequency was applied three times to enhance beta activity. The resulting signal was then rectified and smoothed using a 400-ms moving average to extract beta amplitude. Stimulation was driven by a threshold set at the 75th percentile of the amplitude distribution. During the aDBS procedure, the online beta amplitude was calculated using the same method. Stimulation was initiated when the amplitude exceeded the predefined threshold and was turned off once it fell below that threshold. A minimum stimulation duration of 400 ms was enforced.
3. Proposed algorithm – SMARTA+
The SMARTA+ algorithm builds on the previously proposed SMARTA framework (Shrinkage and Manifold-based Artifact Removal using Template Adaptation) [24]. SMARTA+ introduces two key enhancements. First, to reduce computational burden, the original k-nearest neighbors (KNN) search is replaced with an ANN algorithm, and dimensionality is reduced via wavelet transform. These changes enable the construction of a larger and more diverse stimulus artifact library from available recordings, improving the accuracy of artifact removal. Second, a curve-fitting procedure is incorporated to better handle transient DC artifacts.
Figure 2. Flowchart of SMARTA+.
Fig. 2 is the flowchart of SMARTA+, and the overall workflow is as follows. To address transient DC artifacts, a curve-fitting procedure is applied (Section 3.1) for all recordings. Prior to artifact removal, a stimulus artifact library is constructed from all available recordings (Section 3.2), potentially including data from other subjects. Then, recover each stimulus artifact applying the following algorithm, which is a modified SMARTA. For each recording we want to eliminate the stimulus artifacts, construct a decision forest from all recordings, where the wavelet transform is applied. For each target stimulus artifact, similar artifacts are retrieved from the library using the constructed decision forest and ANN (Section 3.3), where the leaf nodes at the end of the decision paths form a candidate template pool. Within this pool, the $K$ nearest neighbors are identified, with the optimal $K$ selected via greedy search (Section 3.4). The final artifact template is computed as the median of these neighbors and subtracted from the signal segment. The same curve-fitting step is applied again to further remove residual transient DC artifacts (Section 3.1), yielding the final artifact-free LFP. Below, we detail the algorithm step by step.
Assume the stimulation frequency is $f_{\mathrm{sti}}$, which is $130$ Hz in this study. Assume there are $M$ aDBS recordings that we can leverage and we want to recover the LFP of the $(M+1)$th aDBS recording.
3.1. Transient DC artifact removal
For each recording, the transient DC artifacts were removed prior to the segmentation of stimulus artifacts. Over each stimulation period, the trend that contains transient DC artifact is estimated by the following two-step procedure. The first step is applying a moving-average filter with a window length of 10 ms to eliminate large outliers in the signal and the second phase is fitting a sixth-order polynomial function to the smoothed signal. For non-stimulation periods, the trend was estimated and subtracted using the same procedures, except that a first-order polynomial was used for fitting.
3.2. Stimulus artifact library construction
To construct the stimulus artifact library, all available aDBS recordings were considered, and details are listed in the following two subsections.
3.2.1. Stimulus artifacts detection and segmentation
For the $m$-th recording, where $m=1,\ldots,M+1$, the peak-finding algorithm described in [24] was applied to obtain accurate timing of stimulus artifacts. Before identifying peaks, a third-order high-pass Butterworth filter with a cutoff frequency of 300 Hz was used to remove low-frequency noise. Additionally, a moving-average filter with a window length of 1 s was applied to estimate the trend, which was subsequently subtracted from the signal. The absolute value of the signal was than taken, followed by smoothing using a moving-average filter with a window length of 0.5 ms. A sample point was identified as a peak if it was greater than its neighboring points and exceeded a predefined threshold. This threshold was set as the 95th percentile of all sample points. Furthermore, the interval between detected peaks had to be greater than $1/f_{\mathrm{sti}}-\tau$, where $f_{\mathrm{sti}}$ is the stimulation frequency and $\tau$ is a bias-handling factor. In this study, $\tau$ was set to 0.5 ms. The detected peaks marked the locations of stimulus artifacts for segmentation. For aDBS signals, stimulation periods were further identified by detecting intervals between artifacts that were longer than the minimum period set in the aDBS algorithm.
Before segmentation, second-order IIR notch filters were applied to remove 60-Hz line noise and its harmonics up to 3000 Hz, with a Q factor of 200. Afterwards, segmentation was performed, with each segment containing a single artifact. The $i$-th segment started 1 ms before and ended at $1/f_{\mathrm{sti}}-0.5$ ms after the detected artifact location and is denoted as $x_{i}$. A data matrix was then constructed as
(1) $\operatorname{\boldsymbol{X}}_{m}=\left[x_{m,1}\ x_{m,2}\ \ldots\ x_{m,n_{m}}\right]\in\operatorname{\mathbb{R}}^{p\times n_{m}},$
where $p\in\mathbb{N}$ represents the length of each segment, and $n_{m}$ is the number of segments. $\operatorname{\boldsymbol{X}}_{m}$ is a noisy data matrix, which can be written as $\operatorname{\boldsymbol{X}}_{m}=\boldsymbol{S}_{m}+\boldsymbol{N}_{m}$, where $\boldsymbol{S}_{m}\in\operatorname{\mathbb{R}}^{p\times n_{m}}$ represents the stimulus artifacts that we want to recover from $\operatorname{\boldsymbol{X}}_{m}$, and $\boldsymbol{N}_{m}\in\mathbb{R}^{p\times n_{m}}$ contains LFP, which is viewed as noise.
3.2.2. Prepare data for ANN
Instead of processing all segments in $\operatorname{\boldsymbol{X}}_{m}$ at once, they were divided into smaller groups, each containing 500 segments, and the LFP was cleaned up by eOptShrink. This grouping strategy was implemented to speed up the eOptShrink algorithm. For the sake of self-completeness, eOptShrink algorithm can be found in A. After LFP removal by eOptShrink, each segment underwent wavelet transformation using the Haar wavelet and we obtain the wavelet coefficient matrix:
(2) $\boldsymbol{W}_{m}=\left[w_{m,1}\ w_{m,2}\ \ldots\ w_{m,n_{m}}\right]\in\operatorname{\mathbb{R}}^{q\times n_{m}},$
where $q\in\mathbb{N}$ is the number of wavelet coefficients. The average stimulation amplitude, denoted as $a_{m}:=\frac{1}{n_{m}}\sum_{j=1}^{n_{m}}\max x_{j}$, and the median stimulus artifact, denoted as $s_{m}\in\mathbb{R}^{q}$, were computed as key characteristics of the $m$-th recording for subsequent analysis. Feature selection for applying ANN and KNN was then performed as follows. First, a matrix of all median stimulus artifacts $s_{m}$ was constructed: $S=[s_{1}\ s_{2}\ldots s_{M}]\in\operatorname{\mathbb{R}}^{q\times(M+1)}$. For each $i=1,\ldots,q$, the variance of the $i$th row of $S$ was computed and denoted as $s_{v}(i)$, yielding a variance vector $s_{v}\in\mathbb{R}^{q}$. The entries of $s_{v}$ were then sorted in descending order, and their cumulative sum was computed, resulting in a vector $\tilde{s}_{v}\in\mathbb{R}^{q}$. The index $k$ was defined as the largest integer such that $\tilde{s}_{v}(k)<0.99\tilde{s}_{v}(q)$. The indices of the top $k$ entries in the sorted array were then selected as features for the subsequent ANN and KNN analyses.
3.3. Build decision trees
For the $(M+1)$th aDBS signal, find recordings with stimulation amplitude different from $a_{M+1}$ by less than 10%; that is, all $m$ s.t. $\left|\frac{a_{m}-a_{M+1}}{a_{M+1}}\right|\leq 0.1$ were selected as candidates. Among these candidates, the KNN algorithm was applied to identify $Q$ recordings with closest $s_{m}$ to $s_{M+1}$. If fewer than $Q$ candidates were available, all candidates were included. Collect all stimulus artifact as $\operatorname{\mathbb{G}}_{M+1}=\left[w_{1}\ w_{2}\ \ldots\ w_{n^{\prime}}\right]\in\mathbb{R}^{q\times n^{\prime}}$, where $n^{\prime}$ represents the number of segments in the selected $K$ recordings.
Decision trees were then constructed using $\operatorname{\mathbb{G}}_{M+1}$. To build a tree, two points $w_{i}$, $w_{j}$ were randomly selected to generate a hyperplane, which is determined by the middle point $\frac{w_{i}+w_{j}}{2}$ with the normal direction $\frac{w_{i}-w_{j}}{\|w_{i}-w_{j}\|}$. The data was then divided into two subspaces by the hyperplane. Once one subspace contains fewer than $M^{\prime}$ points, that subspace is not further touched. Otherwise, this process was repeated until all subspaces contained fewer than $M^{\prime}$ elements, and hence a tree structure. The resulting subspaces were termed end nodes of the tree, and the data points they contained were referred to as leaves of the tree. Since the selection of points for hyperplane derivation was random, the results from a single decision tree were unstable and potentially inaccurate. To enhance accuracy, $T$ trees were constructed, forming a decision forest. While a larger $T$ improved accuracy and stability, it also increased computational cost.
3.4. Derive template and recover LFPs
For each artifact segment, the corresponding wavelet coefficients was obtained. These coefficients were then used to locate the corresponding leaves in the decision forest. Specifically, for each node in a decision tree, the subspace corresponding to the wavelet coefficient set was determined by the hyperplane of that node. The set continued traversing the tree until it reached an end node, where the contained leaves were considered its neighbors. After searching through the entire forest, the ANN algorithm identified at most $M\times N$ neighbors, collectively denoted as $\operatorname{\boldsymbol{Y}}$.
Next, the KNN was applied to $\operatorname{\boldsymbol{Y}}$, selecting $K$ closest neighbors. The final template for the $i$th stimulus artifact, denoted as $\hat{\boldsymbol{s}}^{*}_{i}$, was then computed by taking the entry-wise median of the $K$ chosen neighbors. The optimal value of $K$ was determined using a greedy search algorithm that minimized the artifact residual (AR) index [26] defined as
(3) $\begin{split}\mathrm{AR}_{i}=&\left|\mathrm{log}\left(\frac{1}{2}\left[\frac{\mathrm{med}|\hat{Z}_{i}-\mathrm{med}(\hat{Z}_{i})|}{\mathrm{med}|Z_{i}-\mathrm{med}(Z_{i})|}+\frac{\mathrm{med}|Z_{i}-\mathrm{med}(Z_{i})|}{\mathrm{med}|\hat{Z}_{i}-\mathrm{med}(\hat{Z}_{i})|}\right]\right)\right.\\ &\left.\times\frac{1}{2}\left[\frac{\mathrm{Q95}|\hat{Z}_{i}-\mathrm{med}(\hat{Z}_{i})|}{\mathrm{Q95}|Z_{i}-\mathrm{med}(Z_{i})|}+\frac{\mathrm{Q95}|Z_{i}-\mathrm{med}(Z_{i})|}{\mathrm{Q95}|\hat{Z}_{i}-\mathrm{med}(\hat{Z}_{i})|}\right]\right|,\end{split}$
where $\hat{Z}_{i}$ represents the first $t_{1}$ ms and $Z_{i}$ the last $t_{2}$ ms of the estimated LFP signal $\boldsymbol{x}_{i}-\hat{\boldsymbol{s}}_{i}^{*}$ obtained after template subtraction. The functions $\mathrm{med}$ and $\mathrm{Q95}$ denote the median and 95th percentile, respectively. In this study, $t_{1}=3$ ms and $t_{2}=5.2$ ms. The optimal $K$ was selected from the candidate values $K=10,20,\ldots,100$ to minimize the AR index.
The overlap-and-add method was then applied to reconstruct the signal, using a window function defined as
(4) $w(t)=\left\{\begin{array}[]{ccl}\sin^{2}\left(\frac{\pi(t-1)}{2g_{1}}\right)&\mathrm{for}&t\leq g_{1}\\ 1&\mathrm{for}&g_{1}<t\leq g_{2}\\ \cos^{2}\left(\frac{\pi(p-t)}{2g_{2}}\right)&\mathrm{for}&g_{2}<t\leq p\end{array}\right$.,
where $g_{1}$ and $g_{2}$ represent the number of overlapping points with the preceding and following segments, respectively. The final artifact-free LFP signal was obtained by subtracting the sequence of artifact templates from the original signal. Additionally, signal trend removal was performed using the smoothing and polynomial fitting method described in Sec. 3.1. Notably, stimulation periods could be identified prior to artifact removal.
3.5. Performance evaluation
3.5.1. Semi-real aDBS recording generation
To compare different stimulus artifact removal algorithms, semi-real aDBS signals were generated using the set of stimulus artifacts extracted in Sec. 3.2.2 for each aDBS signal, and a decision forest for ANN was constructed as described in Sec. 3.3. After artifact detection, trend removal (with the trend preserved for later use), and segmentation, the semi-real stimulus artifact for each segment was generated following the procedure in Sec. 3.4, with the modification that the number of neighbors $K$ for KNN was fixed at 500. Additionally, instead of taking the median, the final result was computed by averaging all neighbors.
With a large $K$, not all selected neighbors closely resembled the target artifacts, but this approach effectively removed the underlying LFP. The overlap-and-add method was then applied to reconstruct the semi-real artifacts, and the previously extracted trends for both stimulation and non-stimulation periods were reintroduced. As a result, the generated signal retained the stimulus artifacts and the trend of the original aDBS signal while minimizing the presence of underlying LFP activity. Finally, the semi-real aDBS signal was obtained by combining the artifact-only signal with a LFP signal recorded in the absence of stimulation.
3.5.2. Evaluation measures
The effectiveness of artifact removal was quantified using three indices: artifact residual (AR), spectral concentration (SC), and normalized mean square error (NMSE), as described in [24]. The AR for the entire signal was computed by averaging the AR index, $\mathrm{AR}_{i}$, across all segments (i=1,,Ni=1,...,N, where $N$ was the total number of segments). The AR index for each segment, $\mathrm{AR}_{i}$, was defined in Eq. 3, where $Z_{i}$ was formed by concatenating the last $t_{2}$ ms of the (i20)(i-20)th segment to the $(i+20)$th segment.
The SC index, which evaluated signal concentration at a target frequency $f_{c}$, was calculated as
(5) $\mathrm{SC}(f_{c})=\frac{\sum_{f\in F_{1}}P_{\hat{Z}}(f)}{\sum_{f\in F_{2}}P_{\hat{Z}}(f)},$
where $P_{\hat{Z}}$ represents the power spectral density of the estimated signal $\hat{Z}$. The two frequency sets were defined as $F_{1}=\{f|f_{c}-1\leq f\leq f_{c}+1\}$ and $F_{2}=\{f|f_{c}-20\leq f\leq f_{c}+20\}$.
The NMSE was calculated as
(6) $\mathrm{NMSE}=10\mathrm{log}_{10}\frac{\sum_{n}(\hat{Z}(n)-Z(n))^{2}}{\sum_{n}Z(n)^{2}},$
where $\hat{Z}$ and $Z$ were the estimated and the ground-truth LFP signals, respectively.
[…]
with no unit, of a true positive event.
For an aDBS algorithm, two key concerns were: (1) how accurately beta events were detected and (2) how many detected events were incorrect. To carry out the temporal event localization analysis, consider the following metrics. The first one was evaluated using recall, defined as
(7) $\mathrm{Recall}=\frac{\#\mathrm{TP}}{\#\mathrm{TP}+\#\mathrm{FN}}.$
The second concern was assessed using precision, defined as
(8) $\mathrm{Precision}=\frac{\#\mathrm{TP}}{\#\mathrm{TP}+\#\mathrm{FP}}.$
These two indices were further summarized using the F1-score, given by
(9) $\mathrm{F1}=\frac{2\times\mathrm{Recall}\times\mathrm{Precision}}{\mathrm{Recall}+\mathrm{Precision}}.$
Recall quantified the proportion of true events detected, precision measured the proportion of detected events that are true, and the F1-score provided a balanced summary of recall and precision.
3.5.3. Other methods for comparison
To validate the performance of SMARTA+, comparative analyses were conducted against SMARTA and two artifact removal techniques: template subtraction and blanking. SMARTA was implemented following the methodology described in [24], without any modifications. Template subtraction was implemented based on the causal setup described in [7], where templates are generated in real time using only past data. Specifically, for the $i$th segment, ensuring that the estimation was computed as the average of the (ik)(i-k)th to (i1)(i-1)th segments, ensuring that the estimation relies solely on preceding information. Notably, although [7] employed a hybrid approach combining both front-end and back-end modules, only the back-end component of template subtraction could be implemented in this study due to data constraints.
Blanking was also implemented in two forms: pulse blanking and transient blanking. Front-end blanking, as commonly applied in embedded aDBS devices [32], was realized using the sample-and-interpolate technique [14] to perform pulse blanking. Each stimulation artifact pulse was blanked over a 2.5 ms window, with the removed samples interpolated to maintain continuity. Transient blanking, implemented according to the method proposed in [1], was applied at the level of entire stimulation periods rather than individual artifact pulses. Specifically, a 550 ms window following the onset of each stimulation period was blanked to suppress the broadband noise associated with DC transient artifacts.
These four methods, including SMARTA+, SMARTA, template subtraction, and blanking, were evaluated under consistent conditions to assess their effectiveness in suppressing stimulus artifacts and preserving relevant neural signals in aDBS recordings.
4. Results
The SMARTA+ algorithm was implemented in MATLAB 2019b and executed on a machine with an AMD Ryzen 5 3500X 6-core processor, 16 GB RAM, and Windows 10. In this study we have $M=34$. For SMARTA+, the parameters were set to $Q=5$, $M^{\prime}=500$, $T=50$.
4.1. Visualization of different algorithms
SMARTA+ was evaluated in both time and frequency domains and compared with SMARTA, template subtraction, and blanking. As shown in Figure 4, although the stimulus artifact signal (Figure 4E) lacks beta activity (13-35 Hz), trend-induced spectral spreading contaminated the beta band. Moreover, transient DC artifacts introduced beta-like activity at stimulation onsets and offsets (Figure 4D), distorting the beta amplitude in the semi-real signal.
In the time domain, artifacts near the stimulation onset differed from those during stable stimulation (Figure 4BC). As a result, SMARTA struggled to find sufficient matching neighbors within the same signal, leading to suboptimal artifact templates. In contrast, due to the help of ANN, SMARTA+ utilized neighbors from other recordings, allowing for more appropriate template generation. While both methods performed similarly during stable stimulation periods (Figure 4C), as shown in Figure 4B, SMARTA+ (red line) outperformed SMARTA (blue line) at the stimulation onsets since SMARTA+ estimated and removed transient DC artifacts within stimulation periods. In the frequency domain (Figure 4E), both SMARTA and SMARTA+ successfully removed the aliasing peak at 42.8 Hz, but SMARTA+ further reduced low-frequency noise compared to SMARTA since the transient DC artifact is removed and stimulus artifact near onset is better recovered. Regarding beta amplitude estimation, SMARTA failed to fully estimate contamination at stimulation onsets and offsets (Figure 4D). In contrast, SMARTA+ achieved better artifact removal at onsets and successfully removed trends after offsets, allowing the base LFP’s beta amplitude to be recovered more accurately.
Figure 4. Validation of SMARTA+ on a semi-real aDBS signal. (A) Waveform of the generated artifact signal (dashed line), semi-real aDBS signal (black line), base LFP (green line), and artifact-free LFPs obtained using SMARTA (blue line) and SMARTA+ (red line). (B) Zoomed-in view of artifacts at the onset of stimulation. (C) Zoomed-in view of artifacts during stable stimulation. (D) Beta amplitude estimated using the aDBS algorithm. (E) Power spectrum derived using Welch’s method.
Figure 5 presents a comparison using a real aDBS signal, demonstrating results consistent with those observed in the semi-real case. SMARTA struggled to fully remove onset artifacts and DC trends in the time domain, while SMARTA+ effectively eliminated them. In the frequency domain, SMARTA+ achieved greater suppression of low-frequency noise. Beta amplitudes were significantly distorted at the stimulation onsets and offsets; SMARTA partially reduced these effects, but SMARTA+ was able to largely remove them.
Figure 5. Validation of SMARTA+ on a real aDBS signal. (A) Waveform of the real aDBS signal (black line), and artifact-free LFPs obtained using SMARTA (blue line) and SMARTA+ (red line). (B) Zoomed-in view of artifacts at the onset of stimulation. (C) Zoomed-in view of artifacts during stable stimulation. (D) Beta amplitude estimated using the aDBS algorithm. (E) Power spectrum derived using Welch’s method.
4.2. Performance evaluation
SMARTA+, SMARTA, template subtraction, and blanking were evaluated using 34 real and semi-real aDBS recording. A quantitative comparison of these methods, alongside template subtraction and blanking, is presented in Table 1 and Table 2 for real signals and semi-real signals, respectively. The accuracy of the estimated artifact-free signals was assessed using NMSE, and artifact removal effectiveness was evaluated using AR in the time domain and SC in the frequency domain. The AR values presented in Table 1 and Table 2 represent the average across all signal segments, while SC values are averaged over specific frequency points: $f_{c}=129.16\times n$ Hz, $42.8+129.16\times m$ Hz, and $86.36+129.16\times m$ Hz for $n=1,2,3$ and $m=0,1,2,3$. The table also lists the average computation time per segment. Both SMARTA and SMARTA+ substantially outperformed template subtraction and blanking methods in all performance metrics, albeit at the cost of higher computational load. Among these two, SMARTA+ performed better than SMARTA in terms of NMSE, AR and SC, indicating that it can more effectively restore the original LFP signal. A significant advantage of SMARTA+ was its reduced computation time, decreasing the average processing time per segment by more than 90% compared to SMARTA. This improvement enhances its feasibility for real-time aDBS applications.
$\mu$ AR SC $\mathrm{T}_{c}$ (ms)
$(\sigma)$
Raw data 2.4892 0.4681 N/A
(0.2914) (0.1263) N/A
Template 1.5138 0.4276 0.0145
subtraction (0.4135) (0.1082) (0.0009)
Pulse 0.5221 0.3777 0.0230
blanking (0.4837) (0.0887) (0.0007)
Transient 0.5560 0.1467 0.0274
blanking (0.1105) (0.1369) (0.0016)
SMARTA 0.1523 0.0789 181.4394
(0.0540) (0.0126) (14.9233)
SMARTA+ 0.1216 0.0762 9.3409
(0.0392) (0.0117) (0.6506)
Table 1. Performance comparison of artifact removal methods on aDBS signals.
$\mu$ NMSE AR SC $\mathrm{T}_{c}$ (ms)
$(\sigma)$
Raw data 26.7102 2.4688 0.4882 N/A
(6.0883) (0.2806) (0.1499) N/A
Template 18.8519 1.5113 0.4186 0.0157
subtraction (6.3660) (0.4505) (0.1588) (0.0022)
Pulse 14.8655 0.5467 0.4625 0.0249
blanking (5.4402) (0.5073) (0.1042) (0.0070)
Transient 14.9663 0.5581 0.1876 0.0284
blanking (7.1509) (0.1116) (0.1725) (0.0020)
SMARTA -0.8739 0.0738 0.0612 83.9569
(5.2589) (0.0146) (0.0048) (1.3635)
SMARTA+ -10.6868 0.0652 0.0602 7.9830
(2.0356) (0.0089) (0.0055) (0.2417)
Table 2. Performance comparison of artifact removal methods on semi-real aDBS signals. $\mu$: mean; $\sigma$: standard deviation; NMSE: normalized mean square error; $\mathrm{NMSE}_{\beta}$: NMSE of $\beta$ band (13-35Hz); AR: artifact residual; SCSC: spectral concentration; $\mathrm{T}_{c}$: computation time per segment; N/A: not applicable.
To further examine artifact removal performance across frequency bands, NMSE values were computed for alpha (4-8 Hz), beta (13-35 Hz), gamma (60-90 Hz), high-frequency oscillation (HFO; 200-400 Hz), and very high-frequency oscillation (VHFO; 400-1000 Hz and 1000-3000 Hz for VHFO1 and VHFO2, respectively) bands. The results are summarized in Table 3. These values were obtained by applying band-pass filters to the signals before NMSE calculation. In the semi-real aDBS signals, the artifact signal was generated separately by averaging over 500 segments before added to the base LFP signal. This extensive averaging effectively removed most of the underlying beta-band activity from the artifacts, resulting in minimal distortion of beta activity in the raw semi-real data, as reflected by the low beta-band NMSE. The alpha band exhibited high NMSE values in raw signals due to low-frequency drift caused by DC transient artifacts. Only SMARTA+, which incorporates a line-fitting algorithm for transient correction, demonstrated a significant reduction in NMSE in this band. In the gamma band, the aliasing of stimulation harmonics, particularly at 86.36 Hz, contributed to elevated NMSE in raw data. SMARTA+ again yielded the lowest NMSE, outperforming other compared algorithms. The HFO and VHFO bands contained a broader spectrum of artifact-induced distortion, including both harmonics and aliasing. As expected, the NMSE values for the raw semi-real data in this band were the highest. While transient blanking significantly improved HFO-band and VHFO-band NMSE compared to pulse blanking and template subtraction, SMARTA+ achieved the most substantial reduction among all methods. This suggests its superior capacity to suppress both time-domain and frequency-domain components of stimulation artifacts. Overall, SMARTA+ outperformed all other methods across all four frequency bands, supporting its effectiveness in restoring LFP signals over the full spectral range.
$\mu$ $\alpha$ $\beta$ $\gamma$ HFO VHFO1 VHFO2
$(\sigma)$
Raw data 8.8129 -2.8386 6.5825 22.0838 22.0623 28.3007
(6.0202) (6.6031) (5.8542) (7.0670) (5.2973) (5.5852)
Template 8.7394 -3.3086 5.0961 6.7619 9.4645 17.4206
subtraction (5.8646) (6.4528) (5.7970) (6.1430) (5.2593) (5.8010)
Pulse 9.9207 -1.5307 8.5451 19.0937 12.8918 16.5090
blanking (6.2107) (6.8008) (5.1215) (6.2944) (5.6151) (5.3941)
Transient 14.3776 3.6873 6.5831 5.4405 2.7675 6.7541
blanking (7.2699) (6.6184) (6.5790) (7.6582) (5.2580) (6.0722)
SMARTA 7.8823 -4.0546 -0.3614 -7.0830 -10.0250 -10.0163
(6.0174) (6.8566) (5.1834) (5.3489) (4.0238) (5.3279)
SMARTA+ -5.8872 -13.5504 -6.9826 -10.3981 -12.1729 -12.4535
(2.0356) (4.4098) (3.5192) (4.9695) (2.9736) 2.3763
Table 3. Band-specific NMSE values for artifact removal methods across alpha (4-8 Hz), beta (13-35 Hz), gamma (60-90 Hz), HFO (200-400 Hz), VHFO1 (400-1000 Hz), and VHFO2 (1000-3000 Hz).
4.3. Temporal Event Localization Analysis
Table 4 presents the accuracy of the beta event detection algorithm, using the baseline beta amplitude extraction procedure described in Sec. 2.1, following the application of various artifact removal methods. Below, “Raw filtering” denotes the condition in which no artifact suppression techniques was applied before evaluating beta amplitude. Transient blanking [1] was intended to avoid artifact-induced self-triggering by suppressing beta detection for a fixed duration following stimulation onset. However, because it simply omitted the first 550 ms of each stimulation period, it did not contribute to accurate beta event detection and could not be meaningfully assessed. As such, transient blanking was excluded from direct performance comparisons.
The performance metrics were averaged across 34 semi-real aDBS signals. To further investigate the efficacy of artifact removal at the initial phase of stimulation, detection performance was also evaluated separately within the first 50 ms of each stimulation period. This allows for a focused comparison of SMARTA+’s effectiveness at mitigating onset-related artifacts.
There are several findings. First, when no artifact removal was applied; that is, raw-filtering, the extracted beta amplitude resulted in high recall but low precision. This pattern was especially pronounced during the onset of stimulation, suggesting that many of the detected beta events were in fact false positives caused by artifacts overlapping with the beta band. The inflated recall, therefore, reflected the detection algorithm’s susceptibility to spurious high-amplitude signals, rather than its ability to identify true neural activity. Second, template subtraction, pulse blanking, and SMARTA all demonstrated improvements upon raw-filtering in detection accuracy by reducing the number of false positives. These method led to a notable increase in precision, thereby improving the F1-score and reducing timing deviation. Among them, SMARTA+ achieved the highest overall performance, particularly at the onset of stimulation. It demonstrated significantly higher precision and F1-score than other approaches, while maintaining a comparable recall. In addition, SMARTA+ notably reduced the timing deviation and increased the overlap ratio in beta event detection, providing more temporally precise alignment between the detected beta activity and the underlying neural signal.
Overall, SMARTA+ outperformed existing algorithms in several critical aspects. It preserved beta activity more effectively while maintaining strong artifact removal performance. Compared to SMARTA, its computational efficiency was significantly improved, making it more suitable for real-time applications. Furthermore, it enhanced the accuracy and timing of beta event detection in aDBS, particularly at stimulation onsets.
Recall (%) Precision (%) F1-score (%) deviation (ms) OR1 (%) OR2 (%)
Onset All Onset All Onset All Onset All Onset All Onset All
Raw filtering 91.32 91.72 57.30 77.25 68.27 83.54 172.37 99.80 71.89 81.30 87.95 90.21
(0.79) (0.38) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) (0.49) (0.30)
Template 93.32 92.02 59.99 77.32 71.16 83.68 177.36 99.56 71.93 80.54 87.40 91.17
subtraction (0.11) (0.52) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) (0.24) (0.02)
Blanking 93.17 91.40 63.23 78.16 73.45 83.98 175.60 101.76 72.48 80.88 86.64 90.33
(0.25) (0.23) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) (0.04) (0.22)
SMARTA 92.73 91.85 63.10 78.82 73.21 84.51 163.59 95.34 74.09 81.66 88.27 90.48
(0.48) (0.41) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) ($\ast$) (0.80) (0.11)
SMARTA+ 91.12 92.18 92.81 93.25 91.81 92.63 59.20 38.65 92.62 94.07 88.72 89.71
Table 4. Temporal event localization accuracy comparison of the aDBS algorithm before and after applying artifact removal methods. The p-value obtained using the paired Wilcoxon signed-rank test between SMARTA+ and another algorithm is shown in the parenthesis, where $\ast$ means $p<10^{-3}$.
5. Discussion
Stimulation-induced artifacts, including stimulus and transient DC components, pose major challenges for aDBS implementation. Effective artifact removal is critical for accurate neural signal detection and reliable feedback control. We proposed SMARTA+, a back-end enhancement of SMARTA [24], aimed at improving computational efficiency and transient DC artifact suppression while maintaining artifact removal performance. Key improvements include replacing the KNN search with ANN and incorporating wavelet-based dimension reduction, reducing computation time to under 10% of that required by SMARTA. SMARTA+ achieved comparable artifact suppression (as measured by AR and SC) and demonstrated improved recovery of clean signals, reflected in lower NMSE over different spectral bands. temporal event localization analysis further confirmed its effectiveness in accurately detecting beta burst events. To our knowledge, this is the first study to evaluate how accurately a stimulus artifact removal algorithm can enhance beta burst detection.
We now discuss how SMARTA+ achieves the encouraging NMSE results. Although the $\mathrm{NMSE}_{\beta}$ for the semi-real aDBS signals was as low as -2.8 shown in Table 3, reflecting the benefits of carefully chosen sampling rates and stimulation frequencies, the high recall but low precision in beta burst detection from raw signals indicates limited effectiveness. Stimulation artifacts might still mislead aDBS algorithms, causing frequent false detections despite the optimized signal design for PD therapy. This issue arises mainly from two factors. First, while the sampling rate and stimulation frequency have been carefully designed to avoid aliasing, the time-varying morphology of stimulus artifacts still introduces spectral leakage via aliasing. To appreciate this statement, consider the following simple model. Suppose the stimulus artifacts are modeled as $f(t)=A(t)\sum_{j\in\mathbb{Z}}s(t-t_{j})$, where $s$ is smooth and compactly supported modeling the morphology of stimulus artifact so that $s(t)=\alpha_{0}+\sum_{k=1}^{\infty}\alpha_{k}\cos(2\pi kt+\beta_{k})$, $\alpha_{0}\in\mathbb{R}$ is the mean value of $s$, $\alpha_{k}\geq 0$ and $\beta_{k}\in[0,2\pi)$, $t_{j}=j/130$ is the timing of stimulation and $A(t)$ is a positive smooth function modeling the strength of stimulus artifact. As discussed in [24], due to the soft matter property of brain, in general the stimulus artifact changes from time to time, particularly the strength. Despite its simple behavior, we know that $\hat{f}(\xi)=\alpha_{0}\hat{A}(\xi)+\frac{1}{2}\sum_{k=1}^{\infty}\alpha_{k}[e^{\beta_{k}}\hat{A}(\xi-130k)+e^{-\beta_{k}}\hat{A}(\xi+130k)]$, where the equality holds in the uniformly. Therefore, when the magnitude $A$ is large, unless $A$ is a constant function or the support of $\hat{A}$ is compact and “very” concentrated around 0, the spectral spreading of $\hat{A}$ could contaminate the beta band via aliasing. Second, abrupt stimulus onsets and transient DC artifacts introduce broadband distortions that overlap with true beta activity, inflating beta amplitude and triggering false detections. This can be specifically seen in the improvement of NMSEα shown in Table 3. Artifact removal using template subtraction, pulse blanking, or SMARTA improves precision and F1-score by reducing these false positives. However, residual transient artifacts, especially at stimulation onset, remain challenging. SMARTA+ addresses this by also targeting onset-related DC transients, resulting in further gains in precision and F1-score, especially during stimulation onset.
A directly related topic is the existing aDBS implementations. On the Summit RC+S system (Medtronic), aDBS is challenged by both stimulus artifacts and DC transients [1]. While front-end blanking effectively removes stimulus artifacts, residual DC transients introduce broadband noise that can degrade beta burst detection and lead to erroneous self-triggering. In [1], this was mitigated by back-end detector blanking with a 550 ms window. However, such a long blanking duration risks omitting relevant beta activity shortly after stimulation, potentially impairing aDBS performance. An alternative approach was explored in [37], where LFPs recorded using the Percept PC system were analyzed. By carefully aligning the stimulation frequency with the sampling rate, stimulus artifacts were minimized within the beta band. However, this study did not implement aDBS or address DC transient artifacts, leaving it unclear whether self-triggering would still occur under aDBS deployment. Also, the spectral leakage via aliasing due to time-varying stimulus artifact morphology is also inevitable in this situation. As a back-end method, SMARTA+ is adaptable to a range of systems, including Summit RC+S and Percept PC, for removing both stimulus and transient artifacts. Given the repeated stimulation epochs in aDBS signals, SMARTA+’s ability to suppress artifacts at each onset supports more reliable beta detection and offers promise for improving adaptive stimulation control.
The amplitude of transient artifacts has been shown to correlate with the rate of change in stimulation amplitude [12]. To mitigate these artifacts, one strategy involves modulating the stimulation amplitude between a predefined upper and lower limit, rather than turning the stimulation fully on or off [40]. This controlled modulation reduces the rate of change and, consequently, diminishes the magnitude of transient artifacts. Nevertheless, despite the reduction in artifact severity, both transient DC artifacts and stimulus artifacts are not entirely eliminated. Furthermore, the effectiveness of this approach is highly dependent on the selection of the lower amplitude limit. In certain cases, this lower bound may approach zero, at which point the resulting transient artifacts can still be substantial. Therefore, even though dual-threshold aDBS algorithms contribute to artifact reduction, the use of SMARTA+ remains beneficial as an additional step to further improve signal quality by suppressing remaining artifacts.
In aDBS applications, stimulation parameters are often restricted to reduce artifact contamination. A typical example is the above-mentioned frequency planning [37] to prevent aliasing peaks from overlapping with the beta band. While cDBS allows more flexibility in tuning stimulation frequency, such options are further limited in aDBS unless artifacts are effectively removed. SMARTA+ lifts this constraint by reliably eliminating stimulus artifacts, thereby allowing broader frequency choices. Another constraint lies in electrode selection. To achieve common-mode rejection, stimulation contact is typically positioned between two recording contacts, limiting viable configurations in clinical practice. Although directional electrodes are increasingly used in DBS to achieve targeted stimulation [38], their use in aDBS has been limited by the necessity of common-mode rejection. Although SMARTA+ was applied to bipolar recordings in this study, it is also effective for unipolar signals. This versatility enables the possibility of integration of directional electrodes into aDBS systems while maintaining robust artifact suppression.
Beta-band power in LFPs is commonly used as the feedback signal for aDBS in Parkinson’s disease. However, relying solely on beta activity may not fully capture the neural dynamics underlying motor and non-motor symptoms [16]. Gamma-band activity, for instance, correlates with movement and motor performance [36, 25], while the alpha band has been linked to cognitive and emotional processing [27], highlighting the potential value of multi-band biomarkers. Yet, incorporating multiple frequency features into aDBS is challenging when stimulation artifacts are not properly addressed. Without effective artifact removal, stimulation frequencies must be carefully chosen to avoid overlap with aliasing peaks across all bands of interest, restricting biomarker flexibility. By suppressing artifact contamination across the full frequency spectrum, SMARTA+ enables more accurate LFP recovery and supports the integration of diverse biomarkers and paves the way for more precise and adaptive aDBS strategies. This will be explored in our future research.
Adaptive DBS has been investigated for conditions beyond PD, using control biomarkers derived from signals other than subthalamic LFPs. In essential tremor, for example, low-frequency oscillations recorded from subdural cortical electrodes over the primary motor cortex and from the ventral intermediate nucleus have been used for aDBS control [29]. While clinical trials for cognitive or psychiatric applications of aDBS are still lacking, neurophysiological studies suggest strong potential in these areas [10]. Such applications may require diverse biomarkers from multiple brain regions and potentially different recording modalities. A key advantage of SMARTA+ is its versatility. It is not restricted to LFPs and can be adapted to various signal types, given appropriate training data. This flexibility positions SMARTA+ as a potential tool for artifact removal across a range of aDBS research settings, enabling the development of more generalizable and disorder-specific aDBS systems. For example, VHFOs (>> 1000 Hz) in EEG signals are relevant to epilepsy [39] and can serve as potential biomarkers. However, these VHFOs can be contaminated by stimulus artifacts if not properly processed. SMARTA+ has a potential to be applied to remove these artifacts, aiding in the development of aDBS systems for the treatment of epilepsy.
This work has several limitations. While SMARTA+ is substantially faster than SMARTA, it remains computationally intensive for real-time use compared with traditional back-end template subtraction or blanking algorithms, particularly when considering battery constraints. Addressing this will require not only advances in energy-efficient hardware but also further algorithmic optimization. Future directions include incorporating randomized algorithms [17, 18] and exploring cloud-based processing. As an algorithm-focused study, we did not explore leveraging hardware properties to further optimize SMARTA+. Such considerations are essential for clinical implementation but lie beyond the current scope and will be addressed in future work. The study population is relatively small and limited to Asian subjects from a single hospital. Larger-scale studies are needed to further validate the proposed algorithm.
6. Conclusion
The SMARTA+ algorithm is proposed as an enhanced and time-efficient extension of SMARTA, with demonstrated effectiveness in removing artifacts in LFP signals recorded during electrical stimulation. For each artifact, an accurate template is constructed by identifying similar segments from both the target signal and a pre-recorded dataset using an approximate nearest neighbors algorithm, guided by a similarity metric based on optimal shrinkage for noise reduction. The underlying LFP signal is then recovered by subtracting this template. Additionally, DC transient artifacts are estimated through a two-stage process involving signal smoothing and polynomial fitting and are subtracted from the signal. Temporal event localization analysis shows that SMARTA+ improves the precision of beta-event detection and reduces the timing deviation of detected events. With a computational cost comparable to the duration of each artifact segment, SMARTA+ represents a promising approach for artifact removal in real-time aDBS algorithms.
7. Ethics statement
This study involving human participants were reviewed and approved by Chang Gung Medical Foundation Institutional Review Board (CGMH-IRB No.202500934B0). The patients/participants provided their written informed consent to participate in this study. Written informed consent was obtained from the individual(s) for the publication of any potentially identifiable images or data included in this article. The research was conducted in accordance with the principles embodied in the Declaration of Helsinki.
8. Funding
This work was supported by the National Science and Technology Council, Taiwan (NSTC112-2314-B-182-042-MY3) and the National Health Research Institutes, Taiwan (NHRI-EX113-11104NI).
Reference
Appendix A eOptShrink
After segmentation, the LFPs in the matrix $\operatorname{\boldsymbol{X}}$ were removed using the extended optimal shrinkage (eOptShrink) algorithm [34]. For the sake of completeness, we summarize eOptShrink here. Without loss of generality, assume $p\leq n$ and denote $\beta=p/n$. Denote the singular value decomposition (SVD) of $\operatorname{\boldsymbol{X}}$ as
(10) $\operatorname{\boldsymbol{X}}=\sum^{p}_{i=1}\tilde{\sigma}_{i}\tilde{\boldsymbol{u}}_{i}\tilde{\boldsymbol{v}}_{i}^{\top},$
where $\tilde{\sigma}_{i}$ is the $i$th singular value, satisfying $\tilde{\sigma}_{1}\geq\tilde{\sigma}_{2}\geq\ldots\geq\tilde{\sigma}_{p}\geq 0$, and $\tilde{\boldsymbol{u}}_{i}$ and $\tilde{\boldsymbol{v}}_{i}$ are the corresponding left and right singular vectors, respectively, with $\|\tilde{\boldsymbol{u}}_{i}\|_{2}=\|\tilde{\boldsymbol{v}}_{i}\|_{2}=1$. Next, the threshold $\hat{\lambda}_{+}$ was computed as
(11) $\hat{\lambda}_{+}:=\tilde{\lambda}_{[n^{1/4}]+1}+\frac{1}{2^{2/3}-1}(\tilde{\lambda}_{[n^{1/4}]+1}-\tilde{\lambda}_{2[n^{1/4}]+1}),$
where $n^{1/4}]$ denotes the closest integer to $n^{1/4}$, and $\tilde{\lambda}_{i}=\tilde{\sigma}_{i}^{2}$ is the $i$th eigenvalue of $\operatorname{\boldsymbol{X}}\operatorname{\boldsymbol{X}}^{\top}$.
For $i=1,\ldots,\hat{r}$, where $\hat{r}$ is the estimated effective rank given by the number of $\tilde{\lambda_{i}}$ greater than $\hat{\lambda}_{+}+n^{-1/3}$; that is,
(12) $\hat{r}=\left|\left\{\tilde{\lambda}_{i}|\tilde{\lambda}_{i}>\hat{\lambda}_{+}+n^{-1/3}\right\}\right|,$
the following value was computed:
(13) $\hat{d}_{i}=\hat{\varphi}_{i}\sqrt{\hat{a}_{1,i}\hat{a}_{2,i}},$
where
(14) $\hat{\varphi}_{i}=\sqrt{\frac{1}{\hat{T}_{i}}},\;\hat{a}_{1,i}=\frac{\hat{m}_{1,i}}{\hat{\varphi}_{i}^{2}\hat{T}_{i}^{{}^{\prime}}},\;\hat{a}_{2,i}=\frac{\hat{m}_{2,i}}{\hat{\varphi}_{i}^{2}\hat{T}_{i}^{{}^{\prime}}},$
(15) $\hat{T}_{i}=\tilde{\lambda}_{i}\hat{m}_{1,i}\hat{m}_{2,i},\;\hat{T}_{i}^{{}^{\prime}}=\hat{m}_{1,i}\hat{m}_{2,i}+\tilde{\lambda}_{i}\hat{m}_{1,i}^{{}^{\prime}}\hat{m}_{2,i}+\tilde{\lambda}_{i}\hat{m}_{1,i}\hat{m}_{2,i}^{{}^{\prime}},$
(16) $\hat{m}_{1,i}:=\frac{1}{p}\left(\sum_{j=1}^{k}\frac{1}{\hat{\lambda}_{j}-\tilde{\lambda}_{i}}+\sum_{j=k+1}^{p}\frac{1}{\tilde{\lambda}_{j}-\tilde{\lambda}_{i}}\right),\;\hat{m}_{2,i}:=\frac{1-\beta}{\tilde{\lambda}_{i}}+\beta\hat{m}_{1,i},$
(17) $\hat{m}_{1,i}^{{}^{\prime}}:=\frac{1}{p}\left(\sum_{j=1}^{k}\frac{1}{(\hat{\lambda}_{j}-\tilde{\lambda}_{i})^{2}}+\sum_{j=k+1}^{p}\frac{1}{(\tilde{\lambda}_{j}-\tilde{\lambda}_{i})^{2}}\right),\;\hat{m}_{2,i}^{{}^{\prime}}:=\frac{1-\beta}{\tilde{\lambda}_{i}^{2}}+\beta\hat{m}_{1,i},$
and $k<p$ is a given parameter ($k=10$ in this study). Additionally,
(18) $\hat{\lambda}_{j}:=\tilde{\lambda}_{k+1}+\frac{1-(\frac{j-1}{k})^{2/3}}{2^{2/3}-1}(\tilde{\lambda}_{2k+1}-\tilde{\lambda}_{k}+1).$
Finally, the LFP-free matrix was obtained as
(19) $\hat{\operatorname{\boldsymbol{S}}}=\sum_{i=1}^{\hat{r}}\hat{d}_{i}\tilde{\boldsymbol{u}}_{i}\tilde{\boldsymbol{v}}_{i}^{\top}=\begin{bmatrix}\hat{\boldsymbol{s}}_{1}&\hat{\boldsymbol{s}}_{2}&\ldots&\hat{\boldsymbol{s}}_{n}\end{bmatrix}$.
To accelerate the eOptShrink algorithm, the randomization technique described in [11] was applied.